# S3-Attention:Attention-Aligned Endogenous Retrieval for Memory-Bounded Long-Context Inference

### v1

@@ D.1 From Information Bottleneck to Feature Matching
We prove this in five steps:
Step 1: Information Bottleneck Framework. The information bottleneck objective is:
[…]
Step 2: Data Processing Inequality. By data processing inequality:
$I(Y;\hat{C}\mid Q)\geq I(Y;\mathcal{F}_{\hat{C}}\mid\mathcal{F}_{Q})$ (11)
Step 3: Disentanglement-Based Decomposition. Under Assumption D.1:
[…]
Step 4: Lower Bound Derivation. Define $\alpha=\min_{f\in\mathcal{F}_{Q}}I(Y;f\mid\mathcal{F}_{Q})>0$. We derive two key lower bounds:
(1) For overlapping features:
[…]
(2) For non-overlapping features:
$\quad\sum_{f\notin\mathcal{F}_{Q}}I(Y;f\mid\mathcal{F}_{Q})\cdot\mathbb{I}(f\in\mathcal{F}_{\hat{C}})\geq-\text{const}.$ (14)
Step 5: Final Inequality. Combining Steps 2-4:
$I(Y;\hat{C}\mid Q)\geq\alpha\cdot|\mathcal{F}_{Q}\cap\mathcal{F}_{\hat{C}}|-\text{const}$ (15)
D.2 IDF Weighting as Compression Regularization
The IDF term in our scoring function (Eq. 7) implicitly penalizes I(C;C^)I($\hat{C}$) in the IB objective:
$\text{IDF}(f)=-\log\frac{|\{t\in C:f\in\mathcal{F}_{t}\}|}{|C|}$ (16)
Effect: Reduces contribution of high-frequency features to $I(Y;\hat{C}\mid Q)$ without reducing I(C;C^)I($\hat{C}$).
D.3 Complexity Analysis
Proposition D.3 (Memory Complexity).
The S3-Attention indexing phase achieves $\mathcal{O}(1)$ GPU memory complexity with respect to context length $L$.
Proof.
Let $B$ be the fixed chunk size ($B=2048$ tokens). Memory usage:
@@ D.3 Complexity Analysis
    Forward pass: $\mathcal{O}(B)$ GPU memory
@@
    SAE encoding: $\mathcal{O}(B\cdot k)$ sparse indices
@@
    Index storage: CPU storage (GPU memory released)
Total GPU memory:
$\mathcal{O}(B\cdot k)=\mathcal{O}(1)$ (17)
Appendix E Why SAE & Why K/Q
@@
Appendix F Extended Qualitative Examples
The following examples visualize where S3-Attention assigns high endogenous scores. These activations should be interpreted as correlational evidence of alignment with internal representations, not as a causal proof that a specific feature “contains” a fact. In some instances, the final answer may be produced partly from the model’s parametric knowledge even when the retrieved spans do not explicitly state the answer; we therefore distinguish “evidence-containing” vs. “cue-like” retrieval cases in our discussion.
@@ Appendix F Extended Qualitative Examples
  • The failure mode or advantage illustrated: For example, recovering a bridge entity missed by RAG, resolving entity-centric questions from internal knowledge when external retrieval fails, or filtering out distractors that share keywords but not causal relevance.
The selected samples cover a diverse range of phenomena, including entity-centric questions (Figures 3(a), 5(a), 8(a)), biochemical and taxonomic reasoning (Figures 3(b), 4(a)), temporal and geographical queries (Figures 7(a), 7(b), 9(a), 9(b)), genre and attribute judgments (Figures 6(b), 10(a)), as well as more abstract relational and historical questions (Figures 4(b), 10(b), 11(a), 11(b)).
@@
Appendix G Analysis of S3 vs. Retrieval-Based Methods (Zero-Shot Setting)
In this section, we analyze the behavior of S3 in comparison with retrieval-based methods under the zero-shot LongBench setting. We emphasize that these results are complementary to the few-shot experiments presented in the main sections, and should be interpreted as an analysis of robustness under minimal prompt supervision, rather than as a replacement for few-shot performance.
Unlike traditional Retrieval-Augmented Generation (RAG), which relies on an external retriever trained independently of the language model (e.g., BGE-M3), S3 can be viewed as an intrinsic compression and retrieval mechanism. It leverages the model’s own attention-derived representations to identify and preserve salient context, without introducing external knowledge beyond what the base model already encodes.
@@
Figure 3: Entity-centric and Biochemical/Taxonomic Reasoning Samples (1/2). Sample 2. Keith Nichol (Entity-centric Question). Query: Prior to playing for Michigan State, Keith Nichol played football for a school located in what city? Description: In this experiment, we compare a traditional Retrieval-Augmented Generation (RAG) baseline with a model equipped with the S3-Attention mechanism on an entity-centric background knowledge question, and the results highlight clear advantages of S3 in semantic retrieval and entity-level knowledge utilization. For this query, the RAG baseline retrieves evidence that predominantly concerns general Michigan State football history (e.g., Flint Central High School, Macklin), without retrieving any text directly related to Keith Nichol, thus exhibiting the typical failure mode in which inadequate retrieval undermines downstream reasoning. By contrast, S3-Attention, under the same conditions, performs activation-based ranking over internal semantic units and assigns a high activation score to the entity “Oklahoma”, despite the absence of explicit external evidence mentioning the target entity. This indicates that S3 goes beyond purely document-level retrieval and leverages semantic-level attention over the model’s pretrained knowledge space to implicitly retrieve relevant entities. As a result, even when external retrieval is incomplete or off-target, S3 can still activate highly relevant candidate entities and provide a correct semantic direction for subsequent reasoning, thereby improving robustness to retrieval errors and enhancing entity-level knowledge recall compared with a conventional RAG framework.
@@
(b) Sample 55. Yoo-hoo and Faygo (Beverage Attributes).
@@
Figure 6: Entity Linking and Attribute Judgment Samples (2/2). Sample 41. Rankin/Bass Production Company. Query: For what type of work is the production company for The Year Without a Santa Claus best known? Description: In Sample 41, the question targets the type of work a production company is best known for. S3’s attention highlights tokens like Santa, Film, and Wonderful, which are closely tied to the holiday-themed animated specials produced by Rankin/Bass. Compared to generic company descriptors (location, legal form, etc.), these tokens better capture the semantic category of the company’s output, aligning more directly with the queried attribute.
Sample 55. Yoo-hoo and Faygo (Beverage Attributes). Query: Are Yoo-hoo and Faygo both carbonated drinks? Description: The core of this query is to determine the carbonated nature of the two beverages. RAG retrieval results focus entirely on the German drink Schorle, with no semantic connection to the target beverages. Due to the inability of exogenous embeddings to capture the attribute information of Yoo-hoo and Faygo, it falls into retrieval failure and cannot respond to the core question at all. In contrast, SAE activates “both” (corresponding to the semantic requirement of “are both”) and “sweet” (a core attribute of beverages), invoking the LLM’s own semantic memory of the two beverages—Yoo-hoo is a non-carbonated chocolate drink and Faygo is a carbonated drink—accurately judging the attribute difference. Its endogenous retrieval does not rely on external embeddings but directly calls the model’s own knowledge base, avoiding the “no relevant results” problem caused by the limitations of RAG’s exogenous retrieval.
(a) Sample 58. Luther: The Calling (Broadcast Year).
(b) Sample 62. Kent Scott (Animated Series End Date).
@@
Figure 7: Temporal and Entity-Work Reasoning Samples (1/2). Sample 58. Luther: The Calling (Broadcast Year). Query: Luther: The Calling is based on the BBC crime drama comprising six episodes that were run in which year? Description: The core of the query is to confirm the broadcast year of the six-episode BBC crime drama that served as the adaptation source. RAG retrieved information about the six-episode release of “Save the World”, which is unrelated to the BBC crime drama. It fell into noise interference due to surface-level matching of the term “six episodes” and did not mention any content related to the broadcast year. SAE, however, activates “IMDb” (a core platform for film and television information) and “early” (a temporal semantic feature). By decoding the LLM’s internal semantic memory of adapted film and television works, it locks in the original drama’s broadcast year, effectively filtering out noise from irrelevant episodes and solving RAG’s “lexical matching noise” problem through deep semantic anchoring.
@@
Sample 62. Kent Scott (Animated Series End Date). Query: When did the animated series Kent Scott wrote end after beginning in September of 2002 on “Nick on CBS”? Description: The query requires clarifying the end date of the animated series. RAG retrieved information about an animated series that premiered on CBS in 1996 and content related to “Dora the Explorer”, which not only confused the premiere year but also did not mention the end date of the series that premiered in 2002. Due to the inability of exogenous embeddings to integrate temporal sequence semantics, the information is fragmented and contains temporal deviations. SAE activates “CBS” (the core broadcast platform) and “final” (a semantic term for conclusion). Through semantic density estimation, it integrates the key information that “the series premiered in 2002 and ended after only one season”, decoding the LLM’s internal temporal chain of “broadcast platform – premiere date – conclusion status”, avoiding RAG’s temporal confusion and information fragmentation, and accurately locking in the end date.
(a) Sample 63. Pamela Adlon (Starred in Tainted).
(b) Sample 66. Deftones (“Radiohead of metal”).
@@
Figure 8: Temporal and Entity-Work Reasoning Samples (2/2). Sample 63. Pamela Adlon (Starred in Tainted). Query: What American actress stars in Tainted? Description: The core of the query is to identify the American actress who starred in Tainted. RAG retrieved information about Pamela Adlon’s identity as an actress, as well as her voice acting, awards, and other career details, but failed to establish a connection between her and Tainted. Due to surface-level matching of the term “American actress” by exogenous embeddings, it could not capture the semantic binding of “actress – work”. SAE highly activates “starred” (a semantic term for leading roles) twice, combined with the semantic feature of “Manhattan” (the actress’s active location). Through the deep connection of “actress – work” within the LLM, it confirms that Pamela Adlon starred in Tainted, making up for RAG’s lack of connection and solving the problem that exogenous retrieval cannot capture specific work bindings.
@@
Sample 66. Deftones (“Radiohead of metal”). Query: Which California band, whose debut album Adrenaline appeared in 1995, has been referred to as “the Radiohead of metal”? Description: The query requires identifying the target California band. RAG mentioned that Adrenaline is the debut studio album of an American alternative metal band and referenced Deftones’ guitarist Stephen Carpenter, but failed to clarify the band’s name. Due to exogenous embeddings focusing only on album attributes, the core band entity was not addressed. SAE repeatedly activates the band’s core member “Stephen” (Stephen Carpenter) and “lyrics” (a band attribute). Through the LLM’s internal semantic connection of “album – band – core member”, it accurately locks in the Deftones band, solving RAG’s lack of core entities and highlighting SAE’s ability to accurately decode semantic chains.
(a) Sample 68. Alsa Mall and Spencer Plaza (Country Location).
(b) Sample 70. Erik Watts’ Father (Birth Year).
@@
Figure 9: Geographical, Genre, and Cultural Reasoning Samples (1/2). Sample 68. Alsa Mall and Spencer Plaza (Country Location). Query: Which country is home to Alsa Mall and Spencer Plaza? Description: The core of the query is to clarify the country where the two malls are located. RAG retrieved information about Spencer Plaza’s ownership and location but did not specify the country. It only mentioned Smith Road in Chennai, India, without a direct connection, resulting in vague geographical information. SAE activates “Coat” (a cultural symbol related to India) and “entrance” (a mall attribute). Through the LLM’s internal geographical memory of well-known malls, it confirms that both malls are located in India, breaking through the vague geographical information caused by RAG’s exogenous retrieval and achieving accurate positioning through deep cultural and geographical semantic binding.
@@
Sample 70. Erik Watts’ Father (Birth Year). Query: When was Erik Watts’ father born? Description: The query requires obtaining the birth year of Erik Watts’ father. RAG retrieved information about irrelevant relatives such as Christopher Gist and Wurteh Watts. Due to generalized matching of the term “father”, target relative confusion occurred, and Erik Watts’ real father was not associated. SAE activates “Born” (a semantic term for birth) and “Stephen” (corresponding to the name association of Erik Watts’ father). By decoding the LLM’s internal semantic features of “person – relative relationship”, it accurately locks in the target relative’s birth year, avoiding RAG’s generalized matching of the term “father” and achieving precise correspondence of relative relationships.
(a) Sample 73. Capital Cities vs. Tweaker (Pop Genre).
(b) Sample 77. Celtic Languages and Territories.
@@
Figure 10: Geographical, Genre, and Cultural Reasoning Samples (2/2). Sample 73. Capital Cities vs. Tweaker (Pop Genre). Query: Which is in the pop genre, Capital Cities or Tweaker? Description: The core of the query is to determine the pop music attribute of the two bands. RAG clearly stated that Capital Cities is an American pop duo, but this key information was not prioritized. Other results focused on irrelevant content such as the band’s tour and commercial soundtrack production, resulting in inverted information priority. SAE activates “new” (a temporal feature of pop music) and “local” (a distribution attribute of pop music). Through semantic density estimation, it prioritizes the key information of “pop duo”, aligning with the LLM’s reasoning demand of “band $\rightarrow$ genre”, optimizing information priority, and avoiding the disconnect between RAG’s score ranking and core needs.
@@
Sample 77. Celtic Languages and Territories. Query: In which six Western European territories have Celtic languages or cultural traits survived? Description: The query requires listing six relevant Western European territories. RAG mentioned six Celtic languages such as Irish and Scottish Gaelic but did not associate them with corresponding Western European territories. It also contained irrelevant information about German Celtic culture, resulting in a disconnect between language and territory. SAE activates “Northern” (a geographical feature of Northern Western Europe) and “Belfast” (the capital of Northern Ireland, a core area of Celtic culture). Through the LLM’s internal semantic connection of “Celtic language – culture – territory”, it accurately locks in six territories including Ireland, Scotland, and Wales, establishing a deep binding between language and geography and solving RAG’s information disconnect problem.
(a) Sample 83. Harold Davis vs. Usain Bolt (Record Event).
(b) Sample 86. Pluralist School and Atomism.
@@
Figure 11: Relational and Historical Reasoning Samples. Sample 83. Harold Davis vs. Usain Bolt (Record Event). Query: In what event was Harold Davis a former record holder, but now is held by Usain Bolt? Description: The core of the query is to determine the specific record-breaking event. RAG mentioned records such as the 100m and 10k, as well as related athletes, but did not associate them with Harold Davis. Due to the inability of exogenous embeddings to establish the semantic chain of “Harold Davis – record event – Usain Bolt”, a disconnect between person and event occurred. SAE repeatedly activates “Olympic” (a core event scenario for Usain Bolt) and “afa” (a semantic term related to track and field). Through the LLM’s internal semantic memory of “former record holder – record event – current holder”, it locks in the 100m sprint event, making up for RAG’s lack of connection between person and event and achieving cross-person semantic chain connection.

### v2

@@ D.6 Complexity Analysis
The implementation processes context in fixed-size chunks of $B=4096$ tokens (Line 22: SCAN_BATCH_SIZE = 4096):
@@
    Forward pass: $\mathcal{O}(B\cdot d)$ GPU memory for key projections
@@
    SAE encoding: $\mathcal{O}(B\cdot k)$ for top-$k$ indices per position
@@
    Index storage: Transferred to CPU after each chunk
Total GPU memory per chunk: $\mathcal{O}(B\cdot d+B\cdot k)=\mathcal{O}(1)$ w.r.t. $L$. ∎
Proposition D.7 (Time Complexity).
Let $L$ be the context length, $k$ the SAE sparsity, and $|\mathcal{F}_{Q}|$ the number of query features.
  • Indexing: $\mathcal{O}(L)$ forward passes, $\mathcal{O}(L\cdot k)$ index insertions
  • Retrieval: $\mathcal{O}(|\mathcal{F}_{Q}|\cdot\bar{n})$ where $\bar{n}$ is the average posting list length
D.7 Connection to Classical IR
Our method can be viewed as a neural extension of classical inverted index retrieval:
Table 7: Correspondence between classical IR and our SAE-based retrieval.
Component Classical IR S3-Attention
Terms Words/n-grams SAE feature indices
TF weighting Term frequency Activation magnitude $a_{f}^{(t)}$
IDF weighting $\log\frac{N}{df}$ $\frac{1}{\log(1+\text{freq})+1}$
Index Inverted index self.indices[layer][f_id]
Scoring BM25 Eq. 9
This perspective provides a more grounded justification: our method inherits the well-established effectiveness of TF-IDF style scoring, applied to learned neural features rather than discrete tokens.
Appendix E IDF Weighting as Adaptive Regularization in Sparse Retrieval
E.0.1 Implementation Details
In our sparse retrieval framework, the relevance score for each context position $t$ is computed as:
$s(t)=\sum_{i\in\mathcal{F}_{q}}\underbrace{a_{i}^{(q)}}_{\text{SAE activation}}\cdot\underbrace{\text{IDF}(f_{i})}_{\text{adaptive weight}}\cdot\mathbb{1}[f_{i}\in\mathcal{F}_{t}]$ (13)
where $\mathcal{F}_{q}$ and $\mathcal{F}_{t}$ denote the active feature sets for query and context position $t$ respectively, and the IDF weight is defined as:
$\text{IDF}(f)=\frac{1}{\log(1+\text{freq}(f))+1}$ (14)
with $\text{freq}(f)$ counting the total occurrences of feature $f$ across all context positions.
E.0.2 Reconciling with the Information Bottleneck Objective
The apparent contradiction arises from a misunderstanding of which information is being compressed. Consider the IB objective:
$\max_{\theta}\;I(Z;Y)-\beta\cdot I(Z;X)$ (15)
where $Z$ represents the retrieved context, $Y$ the answer, and $X$ the full input.
Key Insight:
IDF weighting does not reduce $I(Z;Y)$ uniformly. Instead, it performs selective compression on different types of mutual information:
  • High-frequency features (e.g., common function words, structural patterns): These contribute primarily to $I(Z;X)$ (redundant information about the input) rather than $I(Z;Y)$ (task-relevant information).
  • Low-frequency features (e.g., entity-specific, semantically distinctive): These carry higher task-relevant mutual information $I(Z;Y)$.
The IDF weighting effectively implements:
$I_{\text{effective}}(Z;Y)\approx\sum_{f\in\mathcal{F}}\text{IDF}(f)\cdot I(Z_{f};Y)$ (16)
E.0.3 Why This Maximizes the IB Objective
Proposition.
Under the assumption that feature frequency inversely correlates with task-specificity, IDF weighting increases the ratio $\frac{I(Z;Y)}{I(Z;X)}$.
Justification:
  1. 1.
    High-frequency features have high $I(Z_{f};X)$ but low $I(Z_{f};Y)$ (they appear everywhere, thus non-discriminative for the answer).
  2. 2.
    By down-weighting these features, we reduce $I(Z;X)$ more than $I(Z;Y)$.
  3. 3.
    The implementation also includes a hard threshold (freq > 5000: continue), completely excluding extremely common features that contribute mostly noise.
E.0.4 Corrected Interpretation
The statement “reduces contribution of high-frequency features to I(Z;Y)I(Z;Y)” should be clarified as:
“Reduces the retrieval influence of high-frequency features, which empirically contribute more to $I(Z;X)$ (input redundancy) than to $I(Z;Y)$ (answer utility). This selective down-weighting effectively increases the precision of retrieved context, improving the ratio of task-relevant to task-irrelevant information.”
In essence, IDF weighting acts as an implicit regularizer that approximates the optimal trade-off in the IB objective by leveraging the statistical prior that feature frequency anti-correlates with semantic specificity.
Appendix F Why SAE & Why K/Q
@@
Appendix G Extended Qualitative Examples
The following examples visualize where S3-Attention assigns high endogenous scores. These activations should be interpreted as correlational evidence of alignment with internal representations, not as a causal proof that a specific feature “contains” a fact.
@@ Appendix G Extended Qualitative Examples
  • The failure mode or advantage illustrated: For example, recovering a bridge entity missed by RAG, recovering bridge entities or query-relevant attributes that exogenous retrieval misses, or filtering out distractors that share keywords but not causal relevance.
The selected samples cover a diverse range of phenomena, including entity-centric questions (Figures 5(a), 7(a), 9(b)), biochemical and taxonomic reasoning (Figures 5(b), 6(a)), temporal and geographical queries (Figures 8(b), 9(a), 10(b), 11(a)), genre and attribute judgments (Figures LABEL:fig:sample55, 11(b)), as well as more abstract relational and historical questions (Figures 6(b), 12(a), 12(b), 13(a)).
@@
Appendix H Analysis of S3 vs. Retrieval-Based Methods (Zero-Shot Setting)
In this section, we analyze the behavior of S3 in comparison with retrieval-based methods under the zero-shot LongBench setting. As shown in 8,we emphasize that these results are complementary to the few-shot experiments presented in the main sections, and should be interpreted as an analysis of robustness under minimal prompt supervision, rather than as a replacement for few-shot performance.
Unlike traditional Retrieval-Augmented Generation (RAG), which relies on an external retriever trained independently of the language model (e.g., BGE-M3), S3 can be viewed as an intrinsic compression and retrieval mechanism. It leverages the model’s own attention-derived representations to identify and preserve salient context, without requiring an external retriever or an external corpus beyond the provided input context.
@@
Appendix I Consistency of Chunk-Independent K Projections
A key concern for chunk-independent prefill is whether computing key projections ($\mathbf{K}$) without retaining historical KV states leads to significant deviations from the standard FullKV forward pass. Since our endogenous retrieval mechanism (S3) relies on sparse autoencoder (SAE) features extracted from $\mathbf{K}$, any inconsistency could potentially affect index construction and downstream retrieval behavior.
In this appendix, we quantitatively evaluate the consistency between chunk-independent and FullKV $\mathbf{K}$ projections across layers and chunk sizes, including a stress test at 128,000 tokens.
Setup.
We evaluate Llama-3.1-8B-Instruct on sequences of length up to $\mathbf{128{,}000}$ tokens. We probe four representative layers $\{0,12,16,29\}$ spanning shallow, middle, and deep depths. For a given input sequence, we compare: (i) FullKV, a standard forward pass over the entire sequence with full causal attention; and (ii) Chunk-Independent, processing the same sequence in disjoint chunks of size $B\in\{512,1024,2048,4096\}$, where each chunk is forwarded independently without access to previous KV states. At each selected layer, we extract key projections $\mathbf{K}=W_{K}h$ using forward hooks and evaluate both raw $\mathbf{K}$ vectors and their induced SAE feature representations (SAEs are pretrained and fixed). All experiments are run on cuda:0.
Metrics.
We report complementary consistency metrics: (i) K Cosine Similarity between FullKV and chunk-independent $\mathbf{K}$ vectors, (ii) Relative $\ell_{2}$ Error between corresponding $\mathbf{K}$ vectors (lower is better), (iii) Feature Jaccard Similarity for top-$k$ SAE feature indices per position, and (iv) Retrieval IoU, i.e., intersection-over-union between token positions retrieved using SAE-based inverted indices built from FullKV and chunk-independent representations.
Results.
Table 9 reports results at L=128, [^0] L=128{,}000 tokens (averaged over layers $\{0,12,16,29\}$; layer 0 is trivially identical since no history exists). Overall, chunk-independent prefill produces highly consistent $\mathbf{K}$ projections and SAE features even at 128k context. While deeper layers exhibit small numerical deviations at very small chunk sizes (e.g., Relative $\ell_{2}$ Error up to 0.23 at $B=512$), the induced sparse features remain highly stable (Feature Jaccard $\geq 0.939$), and retrieval decisions are unchanged (Retrieval IoU = 1.0 for all tested chunk sizes and layers). This indicates that any residual numerical differences do not propagate to downstream retrieval behavior in our S3 pipeline.
Table 9: Consistency between FullKV and Chunk-Independent $\mathbf{K}$ Projections at 128k tokens on Llama-3.1-8B-Instruct. We report mean over layers $\{0,12,16,29\}$, with the minimum across layers in parentheses when applicable. Higher is better except for Relative $\ell_{2}$ Error.
Metric B=512 B=1024 B=2048 B=4096
Feature Jaccard Similarity 0.960 (0.939) 0.981 (0.970) 0.998 (0.998) 1.000
K Cosine Similarity 0.964 (0.945) 0.983 (0.975) 0.998 (0.997) 1.000
Retrieval IoU 1.000 1.000 1.000 1.000
Relative $\ell_{2}$ Error $\downarrow$ 0.160 (max 0.230) 0.078 (max 0.111) 0.007 (max 0.010) 0.000
Figure 3: Enlarged view of Figure 2: Endogenous vs. Exogenous Retrieval. Top: RAG (BGE-Small) is distracted by high lexical overlap, ranking a generic biography (Sentence 1) higher than the true answer (Sentence 5). Bottom: S3-Attention (Ours) ignores the distraction, showing sparse activation peaks exclusively at the semantic answer anchor (“The Post”) and its reasoning evidence (“Pentagon Papers”).
Figure 4: Enlarged view of Figure 2: Semantic Attention vs. Lexical Retrieval. Top: RAG (BGE-Small) is distracted by high lexical overlap, ranking a generic biography (Sentence 1) higher than the true answer (Sentence 5). Bottom: S3-Attention (Ours) ignores the distraction, showing sparse activation peaks exclusively at the semantic answer anchor (“The Post”) and its reasoning evidence (“Pentagon Papers”).
@@ Results.
Figure 5: Entity-centric and Biochemical/Taxonomic Reasoning Samples (1/2). Sample 2. Keith Nichol (Entity-centric Question). Query: Prior to playing for Michigan State, Keith Nichol played football for a school located in what city? Description: In this experiment, we compare a traditional Retrieval-Augmented Generation (RAG) baseline with a model equipped with the S3-Attention mechanism on an entity-centric background knowledge question, and the results highlight clear advantages of S3 in semantic retrieval and entity-level knowledge utilization. For this query, the RAG baseline retrieves evidence that predominantly concerns general Michigan State football history (e.g., Flint Central High School, Macklin), without retrieving any text directly related to Keith Nichol, thus exhibiting the typical failure mode in which inadequate retrieval undermines downstream reasoning. By contrast, S3-Attention, under the same conditions, performs activation-based ranking and assigns a high activation score to the entity “Oklahoma”, despite the absence of explicit external evidence mentioning the target entity. This indicates that S3 goes beyond purely document-level retrieval and leverages semantic-level attention to identify relevant entities. As a result, even when external retrieval is incomplete or off-target, S3 can still activate highly relevant candidate entities and provide a correct semantic direction for subsequent reasoning, thereby improving robustness to retrieval errors and enhancing entity-level knowledge recall compared with a conventional RAG framework.
@@
(b) Sample 58. Luther: The Calling (Broadcast Year).
@@
Figure 8: Entity Linking and Temporal Reasoning Samples. Sample 41. Rankin/Bass Production Company. Query: For what type of work is the production company for The Year Without a Santa Claus best known? Description: In Sample 41, the question targets the type of work a production company is best known for. S3’s attention highlights tokens like Santa, Film, and Wonderful, which are closely tied to the holiday-themed animated specials produced by Rankin/Bass. Compared to generic company descriptors (location, legal form, etc.), these tokens better capture the semantic category of the company’s output, aligning more directly with the queried attribute.
@@
Sample 58. Luther: The Calling (Broadcast Year). Query: Luther: The Calling is based on the BBC crime drama comprising six episodes that were run in which year? Description: The core of the query is to confirm the broadcast year of the six-episode BBC crime drama that served as the adaptation source. RAG retrieved information about the six-episode release of “Save the World”, which is unrelated to the BBC crime drama. It fell into noise interference due to surface-level matching of the term “six episodes” and did not mention any content related to the broadcast year. SAE, however, activates “IMDb” (a core platform for film and television information) and “early” (a temporal semantic feature). By prioritizing these key tokens associated with the original drama’s context, it locks in the broadcast year, effectively filtering out noise from irrelevant episodes and solving RAG’s “lexical matching noise” problem through deep semantic anchoring.
(a) Sample 62. Kent Scott (Animated Series End Date).
(b) Sample 63. Pamela Adlon (Starred in Tainted).
@@
Figure 9: Temporal and Entity-Work Reasoning Samples. Sample 62. Kent Scott (Animated Series End Date). Query: When did the animated series Kent Scott wrote end after beginning in September of 2002 on “Nick on CBS”? Description: The query requires clarifying the end date of the animated series. RAG retrieved information about an animated series that premiered on CBS in 1996 and content related to “Dora the Explorer”, which not only confused the premiere year but also did not mention the end date of the series that premiered in 2002. Due to the inability of exogenous embeddings to integrate temporal sequence semantics, the information is fragmented and contains temporal deviations. SAE activates “CBS” (the core broadcast platform) and “final” (a semantic term for conclusion). Through semantic density estimation, it integrates the key information that “the series premiered in 2002 and ended after only one season”, establishing a coherent temporal chain of “broadcast platform – premiere date – conclusion status”, avoiding RAG’s temporal confusion and information fragmentation, and accurately locking in the end date.
@@
Sample 63. Pamela Adlon (Starred in Tainted). Query: What American actress stars in Tainted? Description: The core of the query is to identify the American actress who starred in Tainted. RAG retrieved information about Pamela Adlon’s identity as an actress, as well as her voice acting, awards, and other career details, but failed to establish a connection between her and Tainted. Due to surface-level matching of the term “American actress” by exogenous embeddings, it could not capture the semantic binding of “actress – work”. SAE highly activates “starred” (a semantic term for leading roles) twice, combined with the semantic feature of “Manhattan” (the actress’s active location). Through the deep semantic connection of “actress – work”, it confirms that Pamela Adlon starred in Tainted, making up for RAG’s lack of connection and solving the problem that exogenous retrieval cannot capture specific work bindings.
(a) Sample 66. Deftones (“Radiohead of metal”).
(b) Sample 68. Alsa Mall and Spencer Plaza (Country Location).
@@
Figure 10: Entity-Work and Geographical Reasoning Samples. Sample 66. Deftones (“Radiohead of metal”). Query: Which California band, whose debut album Adrenaline appeared in 1995, has been referred to as “the Radiohead of metal”? Description: The query requires identifying the target California band. RAG mentioned that Adrenaline is the debut studio album of an American alternative metal band and referenced Deftones’ guitarist Stephen Carpenter, but failed to clarify the band’s name. Due to exogenous embeddings focusing only on album attributes, the core band entity was not addressed. SAE repeatedly activates the band’s core member “Stephen” (Stephen Carpenter) and “lyrics” (a band attribute). Through the semantic connection of “album – band – core member”, it accurately locks in the Deftones band, solving RAG’s lack of core entities and highlighting SAE’s ability to accurately decode semantic chains.
@@
Sample 68. Alsa Mall and Spencer Plaza (Country Location). Query: Which country is home to Alsa Mall and Spencer Plaza? Description: The core of the query is to clarify the country where the two malls are located. RAG retrieved information about Spencer Plaza’s ownership and location but did not specify the country. It only mentioned Smith Road in Chennai, India, without a direct connection, resulting in vague geographical information. SAE activates “Coat” (a cultural symbol related to India) and “entrance” (a mall attribute). Through the geographical semantic binding of well-known malls, it confirms that both malls are located in India, breaking through the vague geographical information caused by RAG’s exogenous retrieval and achieving accurate positioning through deep cultural and geographical semantic binding.
(a) Sample 70. Erik Watts’ Father (Birth Year).
(b) Sample 73. Capital Cities vs. Tweaker (Pop Genre).
@@
Figure 11: Relational and Genre Reasoning Samples. Sample 70. Erik Watts’ Father (Birth Year). Query: When was Erik Watts’ father born? Description: The query requires obtaining the birth year of Erik Watts’ father. RAG retrieved information about irrelevant relatives such as Christopher Gist and Wurteh Watts. Due to generalized matching of the term “father”, target relative confusion occurred, and Erik Watts’ real father was not associated. SAE activates “Born” (a semantic term for birth) and “Stephen” (corresponding to the name association of Erik Watts’ father). By analyzing the semantic features of “person – relative relationship”, it accurately locks in the target relative’s birth year, avoiding RAG’s generalized matching of the term “father” and achieving precise correspondence of relative relationships.
@@
Sample 73. Capital Cities vs. Tweaker (Pop Genre). Query: Which is in the pop genre, Capital Cities or Tweaker? Description: The core of the query is to determine the pop music attribute of the two bands. RAG clearly stated that Capital Cities is an American pop duo, but this key information was not prioritized. Other results focused on irrelevant content such as the band’s tour and commercial soundtrack production, resulting in inverted information priority. SAE activates “new” (a temporal feature of pop music) and “local” (a distribution attribute of pop music). Through semantic density estimation, it prioritizes the key information of “pop duo”, aligning with the reasoning demand of “band $\rightarrow$ genre”, optimizing information priority, and avoiding the disconnect between RAG’s score ranking and core needs.
(a) Sample 77. Celtic Languages and Territories.
(b) Sample 83. Harold Davis vs. Usain Bolt (Record Event).
@@
Figure 12: Cultural and Relational Reasoning Samples. Sample 77. Celtic Languages and Territories. Query: In which six Western European territories have Celtic languages or cultural traits survived? Description: The query requires listing six relevant Western European territories. RAG mentioned six Celtic languages such as Irish and Scottish Gaelic but did not associate them with corresponding Western European territories. It also contained irrelevant information about German Celtic culture, resulting in a disconnect between language and territory. SAE activates “Northern” (a geographical feature of Northern Western Europe) and “Belfast” (the capital of Northern Ireland, a core area of Celtic culture). Through the semantic connection of “Celtic language – culture – territory”, it accurately locks in six territories including Ireland, Scotland, and Wales, establishing a deep binding between language and geography and solving RAG’s information disconnect problem.
@@
Sample 83. Harold Davis vs. Usain Bolt (Record Event). Query: In what event was Harold Davis a former record holder, but now is held by Usain Bolt? Description: The core of the query is to determine the specific record-breaking event. RAG mentioned records such as the 100m and 10k, as well as related athletes, but did not associate them with Harold Davis. Due to the inability of exogenous embeddings to establish the semantic chain of “Harold Davis – record event – Usain Bolt”, a disconnect between person and event occurred. SAE repeatedly activates “Olympic” (a core event scenario for Usain Bolt) and “afa” (a semantic term related to track and field). Through the semantic relation of “former record holder – record event – current holder”, it locks in the 100m sprint event, making up for RAG’s lack of connection between person and event and achieving cross-person semantic chain connection.
(a) Sample 86. Pluralist School and Atomism.